# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ 6.6.1 Densification
[…]
@@
6.6.2 Pruning
Removing unimportant Gaussian primitives can ensure efficient representation. In the initial 3DGS framework, opacity was employed as the criterion for determining the significance of a Gaussian. Subsequent research has explored the incorporation of scale as a guiding factor for pruning [74]. However, these approaches primarily focus on individual Gaussian primitives, lacking a comprehensive consideration of the global representation. Therefore, subsequent derivative techniques have addressed this issue.
Importance scores: The volume and the hit count on training views can be introduced along with opacity to jointly determine the global significance score of a Gaussian primitive [24].
@@ 6.6.2 Pruning
where $\gamma(\bm{\Sigma}_{j})$ and $\mathbf{1}(G(\mathbf{X}_{j}),r_{i})$ are volume and the hit count, and $M$ represent the number of training views. Subsequently, Gaussians are ranked according to their global significance scores, and the ones with the lowest scores are pruned. Similar importance scores were improved in other works [200, 201].
@@
Multi-view consistency: Multi-view consistency is a key criterion for determining whether Gaussians need to be pruned. For example, work [126] prunes newly added Gaussians that are not observed by three keyframes within a local keyframe window, while work [131] prunes Gaussians that are invisible in all virtual views but visible in real views.
@@
Distance Metric: Some surface-aware methods often introduce the distance to the surface [118] and SDF values [159], pruning Gaussian primitives that are far from the surface. The distance between Gaussians is also an important metric. GauHuman [62] aims to "merge" Gaussians with small scaling and low KL divergence, as mentioned in Eq. 10.
@@
Learnable control parameter: Introducing a learnable mask based on scale and opacity to determine whether target Gaussian primitives should be removed is also an effective way to prevent 3DGS from becoming overly dense [23].
6.7 Post-Processing
@@
Post-processing strategies for pre-trained Gaussians are important, as they can improve the original efficiency and performance of 3DGS. Common post-processing often improves Gaussian representations through different optimization strategies. This type of work has been discussed in Sec. 6.5.
@@
Representation Conversion: Pre-trained 3DGS can be directly converted to Mesh by introducing Poisson reconstruction [202] on sampled 3D points [156, 160]. Similarly, Gaussian Opacity Fields (GOF) [161] introduces 3D bounding boxes to convert pre-trained 3DGS to a Tetrahedral Grid representation, and then it extracts triangle meshes from it using Binary Search of Level Set. Additionally, LGM [92] first converts pre-trained 3DGS to NeRF representation, and then uses NeRF2Mesh [203] to convert it to Mesh.
@@
Performance and Efficiency: Some works aim to enhance the performance of 3DGS in certain tasks through post-processing, such as multi-scale rendering. Based on pretrained 3DGS, SA-GS [36] introduces a 2D scale-adaptive filter that dynamically adjusts scales based on testing frequency to enhance anti-aliasing performance when zooming out. In terms of efficiency, removing redundant Gaussian primitives from pre-trained 3DGS [29] or introducing a Gaussians caching mechanism [204] can effectively improve rendering efficiency.
6.8 Integration with Other Representations
@@
The convertible nature of 3D representations facilitates the integration of 3DGS with other representations. Such works often leverage the advantages of other representations to improve the original 3DGS.
(a) Extending Triplane to 4D Space [145].
(b) Mapping 3D Gaussian onto a 2D Grid  [25].
Figure 10: Two auxiliary representations of 3DGS.
6.8.1 Point Clouds
Point clouds, as a 3D representation related to 3DGS, are often used to initialize its attributes. Converting point cloud representations to 3DGS can effectively fill in the point clouds’ holes [100, 102], which is typically done after high-precision reconstruction of the point clouds. Conversely, 3DGS can also be converted into point clouds, followed by voxelizing the point cloud into 3D voxels and then projecting them onto 2D BEV grids, which is an important guide in navigation tasks [134].
@@ 6.8.1 Point Clouds
Additionally, anchor points in space can assist 3DGS. These methods use the centers of voxels as anchor points to represent the scene. Each anchor point comprises a local context feature, a scaling factor, and multiple learnable offsets. By decoding other attributes based on these offsets and features, the anchors are transformed into local neural Gaussians. This idea helps mitigate the redundant expansion of 3DGS [26, 39, 159].
6.8.2 Mesh
Meshes have better geometric representation capabilities and can, to some extent, alleviate artifacts or blurry pixels caused by 3DGS [142]. They are still the most widely used 3D representation in downstream tasks [92].
A considerable amount of work has discussed how to convert 3DGS to Mesh, as mentioned in Sec. 5.2. Once converted, they can be further optimized to achieve better geometric and appearance  [64, 75, 76, 163].
@@ 6.8.2 Mesh
Jointly optimizing 3DGS and Mesh is also an optional strategy. 3DGS is suitable for constructing complex geometric structures, while Mesh can be used to reconstruct detailed color appearances on smooth surfaces. Combining the two can enhance reconstruction performance [142].
6.8.3 Triplane
@@
Triplane is notable for its compactness and efficient expressiveness [50], which can be considered an encoded representation. Thanks to these characteristics, triplane is often used in generalization-related tasks. Specifically, a triplane consists of three orthogonal feature planes: the $X$-$Y$ plane, the $Y$-$Z$ plane, and the $X$-$Z$ plane. Features can be obtained by accessing positions in the space, and these features are subsequently decoded to predict Gaussian attributes [60, 93, 50, 51].
Subsequent works [145, 107, 168] attempt to extend the triplane to 4D space ($XYZ$-$T$) by introducing multi-scale HexPlanes [108] or 4D GaussianPlanes [168] to enhance the continuity of 4DGS in the spatiotemporal dimension, as shown in Fig. 10(a).
6.8.4 Grid
@@
Grid is also an efficient representation, as it can access grid corners and interpolate to obtain features or attributes at specific positions. Hash grid [9], as one of the representative methods, has the ability to compress scenes and obtain a more compact and efficient 3DGS [23, 116, 172, 182, 63]. Furthermore, Self-Organizing Gaussian [25] maps the original unstructured 3D Gaussian onto a 2D grid to preserve local spatial relationships, where adjacent Gaussians will have similar attribute values, thereby reducing the memory storage of the scene and maintaining continuity in 3D space, as shown in Fig. 10(b).
6.8.5 Implicit Representation
Implicit representations, benefiting from their representation capability, can be used to mitigate the condition difficulty and surface artifacts of 3DGS [72]. Specifically, introducing NeRF to encode color and opacity can significantly enhance the representation’s adjustability [205]. Moreover, by designing an SDF-to-opacity transformation function [158] or employing mutual geometry supervision [159] to jointly optimize 3DGS and SDF representations, the surface reconstruction performance of 3DGS can be improved.
6.8.6 GaussianVolumes
@@
GaussianVolumes are also utilized for generalizable representations [94], where a volume is composed of a fixed number of 3DGS. This representation maintains the efficiency of Gaussian representations while offering greater manipulability compared to the generalized triplane representation and alleviating the dependence on the accuracy of point cloud predictions.
Discussion: Depending on different needs, various representations can be introduced. However, efficiently converting between different representations is important.
6.9 Guidance by Additional Prior
@@
When dealing with under-determined problems, such as sparse view settings 3.3.1, introducing additional priors is a straightforward method to improve 3DGS performance.
@@
Pre-trained Models: Introducing pre-trained models is an effective paradigm that can guide the optimization through the model’s knowledge. Pre-trained monocular depth models and point cloud prediction models are a common type of prior, where the predicted depth values and positions can be used for the initialization and regularization [52, 54, 102, 131, 100, 55]. Pre-trained 2D image (or 3D and video) generative models are also important in some AIGC-related tasks. They can be used not only for optimization in combination with SDS Loss [80, 196, 103] but also for directly generating (or editing) images for training [100, 91, 102, 109]. Similarly, some works introduce pre-trained image inpainting networks to alleviate difficulties caused by occlusion as well as overlap [77, 100, 102, 162, 171] or super-resolution models for a high level of detail [102, 168] during the generation process. Additionally, pre-trained ControlNet [167] or Large Language Models can also be used to guide 3D generation. The former can enhance geometric consistency under depth guidance [166, 77, 78], while the latter can predict layout maps to guide spatial relationships in multi-object 3D generation scenarios [78]. Notably, certain pre-trained models can endow 3DGS with additional capabilities, such as semantic understanding models, as discussed in Sec. 5.4 and spatial understanding models [131].
@@
More Sensors: Due to the 3D-agnostic nature of 2D images, reconstructing 3DGS can be challenging, especially in large-scale reconstructions such as SLAM and autonomous driving. Therefore, incorporating additional sensors for 3D depth information, including depth sensors [130, 127, 131, 129, 121], LiDAR [112, 133, 116], and optical tactile sensors [55], has the potential to alleviate this issue.
@@
Task-specific Priors: Some reconstruction tasks, such as human reconstruction, target subjects with certain common characteristics. These characteristics, such as template models and Linear Blend Skinning, can be extracted as priors to guide the reconstruction of similar targets. In the reconstruction, animation, and generation of non-rigid objects, many works utilize SMPL [66] and SMAL [206] to provide strong priors for representing the motion and deformation of non-rigid objects like humans [60, 58, 74, 62, 63] and animals [189, 111]. Subsequently, based on the SMPL template, Shell Maps [95] and template meshes are also introduced in combination with 3DGS to address issues of low efficiency in 3DGAN [73] and unclear geometry [65, 64]. Similarly, in head and face reconstruction and animation tasks, some works [68, 70] also use the FLAME model [69] as a prior. Linear Blend Skinning [151] is also employed as prior knowledge to assist in the prediction of 3DGS motion [70, 150]. Additionally, in 3D urban scene reconstruction tasks, HUGS [114] introduces the Unicycle Model to model the motion of dynamic vehicles, thereby making the motion modeling of moving objects smoother.
TABLE II: The Relationships among Challenges, Tasks, and Technological Improvement.
@@
Suboptimal Data Challenges 7.1 Limited Number: Sparse Views  3.3.2, Autonomous Driving Reconstruction  4.3.1, Dynamic 3DGS 5.1 (Monocular Video Part 5.1.2), AIGC 4.2 and Editable 3DGS 5.3. Limited Quality: Slam 4.3.2 and Reconstruction under blurred images (Sec.3.2) or without poses [207, 208] Limited Number: Initialization 6.1, Regularization 6.4, Adaptive Control 6.6 Training Strategies 6.5, and Guidance by Additional Prior 6.9. Limited Quality: Training Strategies 6.5 and Integration with Other Representations (Sec. 6.8)
@@
Generalization Challenges 7.2 Generalization 3.3.1, and Generalization-related tasks in the Human Reconstruction 4.1 and AIGC 4.2 Initialization 6.1, Adaptive Control 6.6, and Integration with Other Representations 6.8
@@
Physics Challenges 7.3 Physical Motion: Dynamic 3DGS 5.1, Physics Simulation 5.5, Animation 5.3.4, Dynamic Human 4.1 and Autonomous Driving Reconstruction  4.3.1. Physical Rendering: Photorealism 3.2 and Physics Simulation 5.5. Physical Motion and Rendering: Attribute Expansion 6.2, Regularization Strategy 6.4, and Guidance by Additional Prior 6.9.
@@
Realness and Efficiency Challenges 7.4 Realness: Photorealism 3.2, Surface Reconstruction 5.2, Semantic Understanding 5.4, some AIGC-related 4.2 and Autonomous Driving 4.3 works. Efficiency: Efficiency 3.1, some works in Autonomous Driving 4.3 and Semantic Understanding 5.4 Realness: Most of the technologies in 6. Efficiency: Attribute Expansion 6.2, Post-Processing 6.7, Adaptive Control 6.6 and Splatting 6.3.
7 Interrelationships and Challenges
As previously mentioned in Sec.3, Sec.4, and Sec.5, there is a substantial amount of research focused on the optimization, application, and extension of 3DGS. However, these discussions often overcomplicate the issue. In this section, we aim to summarize the commonalities across different tasks and propose four core challenges along with their corresponding technical discussions, as shown in Tab.II. Recognizing these common challenges and solutions can simplify research efforts and promote interdisciplinary progress.
7.1 Suboptimal Data Challenges

### v2

@@ VI-F1 Densification
$\text{GDS}:\quad\|\bm{\mu}_{1}-\bm{\mu}_{2}\|^{2}+\text{tr}\left(\bm{\Sigma}_{1}+\bm{\Sigma}_{2}-2(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{2}\bm{\Sigma}_{1}^{-1})^{1/2}\right),$
$\text{KL}:\text{\small\small$\displaystyle\frac{1}{2}\left(\text{tr}(\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1})+\ln\frac{\det\bm{\Sigma}_{1}}{\det\bm{\Sigma}_{2}}+(\bm{\mu}_{2}-\bm{\mu}_{1})^{T}\bm{\Sigma}_{2}^{-1}(\bm{\mu}_{2}-\bm{\mu}_{1})-3\right).$}$ (10)
@@
VI-F2 Pruning
Removing unimportant Gaussian primitives can ensure efficient representation. In the initial 3DGS framework, opacity was employed as the criterion for determining the significance of a Gaussian. Subsequent research has explored the incorporation of scale as a guiding factor or distractor masks for pruning [93, 49]. However, these approaches primarily focus on individual Gaussian primitives, lacking a comprehensive consideration of the global representation. Therefore, subsequent derivative techniques are discussed.
Importance scores: The volume and hit count on training views, along with opacity, can be used to jointly determine the global significance score of a Gaussian primitive [15].
@@ VI-F2 Pruning
where $\gamma(\bm{\Sigma}_{j})$ and $\mathbf{1}(G(\mathbf{X}_{j}),r_{i})$ are volume and hit count, and $M$ represents the number of training views. Subsequently, Gaussians are ranked according to their global scores, and the ones with the lowest scores are pruned. Similar importance scores were improved in other works [255, 256].
@@
Multi-view consistency: Multi-view consistency is a key criterion for determining whether Gaussians need to be pruned. For example, [238] prunes newly added Gaussians that are not observed by three keyframes within a local keyframe window, while [162] prunes Gaussians that are invisible in all virtual views but visible in real views.
@@
Distance Metric: Surface-aware methods often use distance to the surface [151] and SDF values [195] to prune Gaussian primitives far from the surface. The distance between Gaussians is also a key metric [179]. GauHuman [69] aims to merge Gaussians with small scale and low KL divergence, as mentioned in Eq. 10.
@@
Learnable control parameter: Introducing a learnable mask based on scale and opacity to determine whether Gaussian primitives should be removed effectively prevents 3DGS from becoming overly dense [14].
Others: CoR-GS [63] aims to leverage mismatched regions between two 3DGS models, trained in parallel under identical conditions, as guidance for pruning.
Discussion: Adaptive Control strategies play a pivotal role in enhancing rendering fidelity and computational efficiency. However, excessive densification or pruning can adversely affect both the efficiency and performance of 3D Gaussian Splatting. Therefore, it is crucial to examine and establish an optimal balance between these two strategic approaches.
VII Other Technical Discussions
Several existing works have explored the augmentation of the original 3DGS pipeline through the integration of supplementary components, including post-processing modules and additional information as well as representations. These architectural enhancements typically yield improved performance metrics and facilitate optimization under ill-posed conditions.
VII-A Post-Processing
@@
Post-processing strategies for pre-trained Gaussians are important, as they can improve the original efficiency and performance. Common post-processing often enhances Gaussian representations through various optimization strategies. This type of work has been discussed in Sec. VI-E.
@@
Representation Conversion: Pre-trained 3DGS can be converted to Mesh using Poisson reconstruction [257] on sampled 3D points [192, 197]. Similarly, GOF [198] uses 3D bounding boxes to convert 3DGS to a Tetrahedral Grid, then extracts meshes using Binary Search of Level Set. Additionally, LGM [115] converts Pre-trained 3DGS to NeRF, then uses NeRF2Mesh [258] for Mesh conversion.
@@
Performance and Efficiency: Some works enhance 3DGS performance in specific tasks through post-processing, such as multi-scale rendering. SA-GS [32] introduces a 2D scale-adaptive filter to dynamically adjust scales based on rendering frequency, enhancing anti-aliasing when zooming out. For efficiency, removing redundant Gaussian primitives from pre-trained 3DGS [23] or introducing a Gaussian caching mechanism [259] can improve rendering efficiency.
VII-B Integration with Other Representations
@@
The convertible nature of 3D representations facilitates the integration of 3DGS with other representations, leveraging their advantages to improve the original 3DGS.
VII-B1 Point Clouds
@@
Point clouds, as a 3D representation related to 3DGS, are often used to initialize positions. Converting point clouds to 3DGS can effectively fill holes [129, 130] or improve reconstruction details [119], typically after high-precision reconstruction. Conversely, 3DGS can be converted into point clouds, voxelized into 3D voxels, and projected onto 2D BEV grids, which guide navigation tasks [164]. Additionally, anchor points in space can assist 3DGS. These methods use voxel centers as anchor points to represent the scene. Each anchor point comprises a local context feature, a scaling factor, and multiple learnable offsets. By decoding other attributes based on these offsets and features, the anchors transform into local neural Gaussians, which helps mitigate redundant expansion [16, 36, 195].
VII-B2 Mesh
@@
Meshes have better geometric representation capabilities and can, to some extent, alleviate artifacts or blurry pixels caused by 3DGS [173]. They are still the most widely used 3D representation in downstream tasks [115]. Much work has discussed converting 3DGS to Mesh, as mentioned in Sec. V-B. Once converted, they can be optimized for better geometry and appearance [71, 95, 96, 202]. Jointly optimizing 3DGS and Mesh is also an optional strategy. 3DGS is suitable for constructing complex geometric structures, while Mesh can be used to reconstruct detailed color appearances on smooth surfaces. Combining the two can enhance reconstruction performance [173] and large-scale deformation control [237].
VII-B3 Triplane
@@
Triplane, known for its compactness and efficient expressiveness [51], is often used in generalization tasks. It consists of three orthogonal feature planes: $X$-$Y$, $Y$-$Z$, and $X$-$Z$. Features can be obtained by querying positions in the space, and subsequently decoded to predict Gaussian attributes [67, 114, 51, 52]. Recent works [176, 139, 208, 179] extend triplane to 4D space ($XYZ$-$T$) using multi-scale HexPlanes [143] or 4D GaussianPlanes [208] to enhance 4DGS continuity in the spatiotemporal dimension.
VII-B4 Grid
@@
Grid is also an efficient representation, as it can access grid corners and interpolate to obtain features or attributes at specific positions. Hash grid [171], a representative method, can compress scenes and achieve a more compact and efficient 3DGS [14, 150, 213, 228, 70, 118, 19, 120]. Furthermore, Self-Organizing Gaussian [18] maps unstructured 3D Gaussians onto a 2D grid to preserve local spatial relationships, where adjacent Gaussians have similar attribute values, reducing memory storage and maintaining continuity in 3D space.
@@
Particularly, GaussianVolumes are also used for generalizable representations [118], where a volume is composed of a fixed number of 3DGS. This maintains the efficiency of 3DGS and offers greater manipulability compared to triplane.
VII-B5 Implicit Representation
Implicit representations, benefiting from their representational capability, can be used to mitigate the condition difficulty and surface artifacts of 3DGS [91]. Specifically, introducing NeRF to encode color and opacity can significantly enhance the representation’s adjustability [260]. Moreover, by designing an SDF-to-opacity transformation function [194] or employing mutual geometry supervision [195] to jointly optimize 3DGS and SDF representations, the surface reconstruction performance of 3DGS can be improved.
Discussion: Given the inherently unstructured characteristics of 3DGS, the incorporation of structured representations emerges as a viable prior, particularly advantageous for tasks such as human body reconstruction, facilitating both cross-representation transformations and geometric reconstruction, thereby enabling enhanced performance in downstream applications. However, due to certain limitations in representational characteristics, this approach may be accompanied by some degradation in rendering performance.
VII-C Guidance by Additional Prior
@@
When dealing with under-determined problems, such as sparse view settings III-C, introducing additional priors is a straightforward method to improve 3DGS performance.
@@
Pre-trained Models: Introducing pre-trained models is an effective paradigm that can guide the optimization through the model’s knowledge. Pre-trained monocular depth models and point cloud prediction models are a common type of priors, where the predicted depth values and positions can be used for the initialization and regularization [62, 56, 57, 130, 162, 129, 58]. Pre-trained 2D image (or 3D and video) generative models are also important in some AIGC-related tasks. They can be used not only for optimization in combination with SDS Loss [100, 247, 133] but also for directly generating (or editing) images for training [129, 109, 130, 140, 61]. Similarly, some works introduce pre-trained image inpainting networks to alleviate difficulties caused by occlusion as well as overlap [97, 129, 130, 201, 211] or super-resolution models for a high level of detail [130, 208] during the generation process. Additionally, pre-trained ControlNet [207] or Large Language Models can also be used to guide 3D generation. The former can enhance geometric consistency under depth guidance [205, 97, 98], while the latter can predict layout maps to guide spatial relationships in multi-object 3D generation scenarios [98]. Notably, certain pre-trained models can endow 3DGS with additional capabilities, such as semantic understanding models, as discussed in Sec. V-E and spatial understanding models [162].
@@
More Sensors: Due to the 3D-agnostic nature of 2D images, reconstructing 3DGS can be challenging, especially in large-scale reconstructions such as SLAM and autonomous driving. Therefore, incorporating additional sensors for 3D depth information, including depth sensors [161, 156, 162, 250, 154], audio [87, 88, 89], LiDAR [50, 146, 163, 150, 149], and optical tactile sensors [58], has the potential to alleviate this issue.
@@
Task-specific Priors: Some reconstruction tasks, such as human reconstruction, target subjects with certain common characteristics. These characteristics, such as template models and Linear Blend Skinning, can be extracted as priors to guide the reconstruction of similar targets. In the reconstruction, animation, and generation of non-rigid objects, many works utilize SMPL [74] and SMAL [261] to provide strong priors for representing the motion and deformation of non-rigid objects like humans [67, 65, 93, 69, 70] and animals [239, 144]. Subsequently, based on the SMPL template, Shell Maps [262] and template meshes are also introduced in combination with 3DGS to address issues of low efficiency in 3DGAN [92, 94] and unclear geometry [72, 71]. Similarly, in head and face reconstruction and animation tasks, some works [77, 82] also use the FLAME model [81] as a prior. Linear Blend Skinning [263] is also employed as prior knowledge to assist in the prediction of 3DGS motion [82, 184]. Additionally, in 3D urban scene reconstruction tasks, HUGS [148] introduces the Unicycle Model to model the motion of vehicles, thereby making the motion modeling of moving objects smoother.
Discussion: Accessible auxiliary information has the capability to enhance the performance of 3DGS across numerous tasks, serving as prior knowledge to facilitate spatial comprehension, particularly in inherently ill-posed problems. Although certain priors or sensors may lead to increased computational overhead and costs, they have the capability to significantly enhance the representational capacity of 3DGS .
TABLE V: The Relationships among Challenges, Tasks, and Technological Improvement, where the first column represents the core challenges, while the second and third columns denote the related downstream tasks and the corresponding technological advancements associated with these challenges, respectively
@@
Suboptimal Data   Challenges VIII-B Limited Number: Sparse Views  III-D, Autonomous Driving Reconstruction  IV-C1, Dynamic 3DGS V-A (Monocular Video Part V-A2), AIGC IV-B and Editable 3DGS V-C. Limited Quality: Slam IV-C2 and Reconstruction under blurred images (Sec.III-B) or without poses [264, 265] Limited Number: Initialization VI-A, Regularization VI-D, Adaptive Control VI-F Training Strategies VI-E, and Guidance by Additional Prior VII-C. Limited Quality: Training Strategies VI-E and Integration with Other Representations (Sec. VII-B)
@@
Generalization Challenges VIII-C Generalization III-C, and Generalization-related tasks in the Human Reconstruction IV-A and AIGC IV-B Initialization VI-A, Adaptive Control VI-F, and Integration with Other Representations VII-B
@@
Physics Challenges VIII-D Physical Motion: Dynamic 3DGS V-A, Physics Simulation V-F, Animation V-C4, Dynamic Human IV-A and Autonomous Driving Reconstruction  IV-C1. Physical Rendering: Photorealism III-B and Physics Simulation V-F. Physical Motion and Rendering: Attribute Expansion VI-B, Regularization Strategy VI-D, and Guidance by Additional Prior VII-C.
@@
Realness and Efficiency Challenges VIII-E Realness: Photorealism III-B, Surface Reconstruction V-B, Semantic Understanding V-E, some AIGC-related IV-B and Autonomous Driving IV-C works. Efficiency: Efficiency III-A, some works in Autonomous Driving IV-C and Semantic Understanding V-E Realness: Most of the technologies in VI. Efficiency: Attribute Expansion VI-B, Post-Processing VII-A, Adaptive Control VI-F and Splatting VI-C.
VIII Challenges and Opportunities
The preceding discussion indicates that various 3DGS-related tasks share similar technical approaches, which stems from common challenges across different tasks. To provide readers with a deeper understanding of this phenomenon, this section examines the commonalities among different tasks, summarizes four core challenges as well as their corresponding technical solutions, and outlines future opportunities.
VIII-A Interrelationships
We have extensively discussed various 3DGS-related tasks in Sec.III, Sec.IV, and Sec. V, revealing common challenges and techniques across these tasks. As illustrated in Tab.V, we categorize existing tasks according to four core challenges, demonstrating that solutions from different tasks can be mutually instructive. Furthermore, there are some interrelationships between different tasks that have not been mentioned. For instance, Surface Reconstruction techniques (Sec.V-B) are often referenced in the context of Editable 3DGS (Sec.V-C), etc. We anticipate that this analysis will offer valuable insights for future research endeavors in related tasks.
VIII-B Suboptimal Data